OpenStudy (anonymous):

What is the value of f(6) for the function f(x) = 3x – 5?

OpenStudy (anonymous):

Also what is: What is the value of g[f(3)] for the functions f(x) = 3x – 5 and g(x) = x + 1? PLEASE HELP!

OpenStudy (anonymous):

f(6)=3(6)-5=18-5=13

OpenStudy (anonymous):

f(x)=3x-5 ,g(x)=x+1 ---->f(3)=3(3)-5=4 ---->g(4)=4+1=5

OpenStudy (anonymous):

got it??:D

OpenStudy (anonymous):

Yes! Thanks so very much :D :D